### v2

@@ 5.2 Riemannian gradient and Hessian operator on $\mathcal{C}_{p_{1},p_{2}}^{++}$
Endow $\mathcal{C}_{p_{1},p_{2}}^{++}$ with the Euclidean metric $g^{E}$. By Proposition 6, the form of the tangent space $T_{C}\mathcal{C}_{p_{1},p_{2}}^{++}$ does not depend on $C\in\mathcal{C}_{p_{1},p_{2}}^{++}$. The same holds for the form of the metric, i.e., $g_{C}^{E}(U,V)=\textsf{tr}\left(U^{\top}V\right)$ for $U,V\in T_{C}\mathcal{C}_{p_{1},p_{2}}^{++}$. Thus, letting $\mathcal{W}:=T_{C}\mathcal{C}_{p_{1},p_{2}}^{++}$, which does not depend on $C$, $\mathcal{W}$ is a linear subspace of $\mathcal{S}_{p}$. Also, with any fixed basis of $\mathcal{W}$ as a coordinate on each tangent space, we have that the Christoffel symbols (see $(5.10)$ of [lee2018]) vanish on that coordinate. Thus, $(\mathcal{C}_{p_{1},p_{2}}^{++},g^{E})$ has a zero-sectional curvature, and so $(\mathcal{C}_{p_{1},p_{2}}^{++},g^{E})$ is flat ([lee2018], Theorem $7.10$). Now we derive the Riemannian gradient and Hessian operator on $(\mathcal{C}_{p_{1},p_{2}}^{++},g^{E})$. For a scalar-valued smooth map $f$ on $\mathcal{C}_{p_{1},p_{2}}^{++}$, denote the Euclidean derivative and Hessian operator of $f$ by $\nabla f(C)$ and $\nabla^{2}f(C)[V]$ for $C\in\mathcal{C}_{p_{1},p_{2}}^{++}$ and $V\in T_{C}\mathcal{C}_{p_{1},p_{2}}^{++}$.
@@
(25) $\mathcal{G}(V):=V-\frac{1}{p_{2}}(I_{p_{2}}\otimes\textsf{tr}_{1}(V))-\frac{1}{p_{1}}(\textsf{tr}_{2}(V)\otimes I_{p_{1}})+\frac{\textsf{tr}(V)}{p}I_{p}$
@@
Proof 5.6.
See Appendix A.6 for the proof.
Proposition 13.
@@
Proof 5.7.
See Appendix A.6 for the proof.
@@
We derive the Riemannian gradient and Hessian operator on $(\mathcal{C}_{p_{1},p_{2},r},g^{E})$ and then deduce those on $(\mathcal{C}_{p_{1},p_{2},r}/\mathcal{O}_{r},g^{E,0})$ via quotient geometry. Here $g^{E,0}$ is the quotient metric induced by $g^{E}$. Throughout this section, we denote the Moore-Penrose pseudoinverse of a matrix $M$ by $M^{\dagger}$ (see p.50–51 of [rao1971]). We establish the Riemannian gradient and Hessian operator on $(\mathcal{C}_{p_{1},p_{2},r},g^{E})$ as follows.
@@
Proof 5.9.
See Appendix A.7 for the proof.
Proposition 14.
@@
Proof 5.10.
See Appendix A.7 for the proof.
The Riemannian gradient and Hessian operator on $(\mathcal{C}_{p_{1},p_{2},r}/\mathcal{O}_{r},g^{E,0})$ can be derived using the results of [absil2008, massart2020] (see Section $5.1$ and $5.6$ of [chen2025] for example). For $A\in\mathcal{C}_{p_{1},p_{2},r}$, the vertical and horizontal spaces at $A$ are given by
@@
Note that $T_{A}\mathcal{C}_{p_{1},p_{2},r}=\mathcal{V}_{A}\oplus\mathcal{H}_{A}$. Then we introduce the operators on $T_{A}\mathcal{C}_{p_{1},p_{2},r}$ by [massart2020] as
(26) $P_{A}^{v}(W):=A\mathbf{T}_{A^{\top}A}^{-1}(2\textsf{skew}(A^{\top}W)),\quad P_{A}^{h}(W):=W-P_{A}^{v}(W).$
for $W\in T_{A}\mathcal{C}_{p_{1},p_{2},r}$. Here the operator $\mathbf{T}_{E}^{-1}(\cdot)$ is the inverse of the map $\mathbf{T}_{E}:Y\in\real^{r\times r}\rightarrow YE+EY\in\real^{r\times r}$. If $E\in\mathcal{S}_{r}^{++}$, $\mathbf{T}_{E}$ is invertible and the value of its inverse $\mathbf{T}_{E}^{-1}(V)$ for given $V$ is a unique solution to the Sylvester equation $YE+EY=V$ ([massart2020], Lemma A.10). Also, by Section $5.3.4$ of [absil2008], the Riemannian connection $\nabla$ on $\mathcal{C}_{p_{1},p_{2},r}/\mathcal{O}_{r}$ satisfies that
@@
for any $A\in\mathcal{C}_{p_{1},p_{2},r}$, vector fields $\eta,\xi$ on $\mathcal{C}_{p_{1},p_{2},r}/\mathcal{O}_{r}$ and the operator $P_{A}^{h}$ defined in (26). Note that $\eta^{\#}$ and $\xi^{\#}$ are horizontal lifts of $\eta$ and $\xi$, respectively. As a direct consequence of Section $3.6.2$ and $5$ of [absil2008] and Proposition A.14 of [massart2020], we have the following results.
Proposition 15.
@@ 6 Partial isotropy core shrinkage estimator
Using the geometry of $\mathcal{C}_{p_{1},p_{2},r}$, we shall propose a shrinkage estimator that shrinks the low-dimensional core toward the trivial core, $I_{p}$. Suppose $Y_{1},\ldots,Y_{n}\overset{i.i.d.}{\sim}N_{p_{1}\times p_{2}}(0,\Sigma)$ and let $K^{1/2}CK^{1/2,\top}$ be a KCD of $\Sigma$. In the estimation of $\Sigma$, note that the dimension of the space where $K=k(\Sigma)$ is living is $O(p_{1}^{2}+p_{2}^{2})=o(p^{2})$, whereas that of the space where $C=c(\Sigma)$ is living is $O(p^{2})$ by Theorem 5. Thus, the difficulty of the estimation arises mainly from estimating $C$, particularly in a high-dimensional regime where $p>n$.
@@
As discussed in Section 1 and 2.1, one remedy is to introduce a low-dimensional structure to $C$. In particular, we assume that $C$ exhibits a partial-isotropy structure. By Proposition 1, $C=(1-\lambda)AA^{\top}+\lambda I_{p}$ for some $\lambda\in(0,1)$ and $A\in\mathcal{C}_{p_{1},p_{2},r}$, with $r>p_{1}/p_{2}+p_{2}/p_{1}$. Thus, $\Sigma=K^{1/2}((1-\lambda)AA^{\top}+\lambda I_{p})K^{1/2,\top}$. This implies that $\lambda$ denotes the shrinkage amount of the low-dimensional covariance $K^{1/2}AA^{\top}K^{1/2,\top}$ toward the separable part $K:=K^{1/2}K^{1/2,\top}$. Hence, $\lambda$ quantifies an effective departure from the separability assumption on $\Sigma$, i.e., $\Sigma=K$, where the correlation structure of $\Sigma$ may be too simplified as $p$ grows. We consequently propose the partial-isotropy core covariance model as
(27) $Y_{1},\ldots,Y_{n}\overset{i.i.d.}{\sim}N_{p_{1}\times p_{2}}(0,K^{1/2}((1-\lambda)AA^{\top}+\lambda I_{p})K^{1/2,\top}).$
@@
We shall propose a partial-isotropy core shrinkage estimator (PICSE) by $\hat{K}^{1/2}((1-\hat{\lambda})\hat{A}\hat{A}^{\top}+\hat{\lambda})\hat{K}^{1/2,\top}$, where $\hat{\theta}$ is a MLE of the parameter $\theta$. We consider boths square roots $K^{1/2}\in\mathcal{S}_{p_{1},p_{2}}^{++}$ or $\mathcal{L}_{p_{1},p_{2}}^{++}$. To avoid the ambiguity due to a constant factor in identifying the factors of $K^{1/2}:=K_{2}\otimes K_{1}$, we parameterize $K^{1/2}$ by $K^{1/2}:=\nu(\bar{K}_{2}\otimes\bar{K}_{1})$, where $\bar{K}_{i}\in\mathbb{P}(\mathcal{S}_{p_{i}}^{++})$ or $\mathbb{P}(\mathcal{L}_{p_{i}}^{++})$, and $\nu>0$. Then the negative log-likelihood is given by
(28) $\begin{split}\ell(\bar{K}_{1},\bar{K}_{2},\nu,A,\lambda)&:=\textsf{tr}\left(\bar{K}^{-1/2}S\bar{K}^{-1/2,\top}((1-\lambda)AA^{\top}+\lambda I_{p})^{-1}\right)/\nu^{2}\\ &+\log|(1-\lambda)AA^{\top}+\lambda I_{p}|+2p\log\nu.\end{split}$
@@
(29) $\hat{\Sigma}_{\texttt{PICSE}}:=\hat{\nu}^{2}(\hat{\bar{K}}_{2}\otimes\hat{\bar{K}}_{1})((1-\hat{\lambda})\hat{A}\hat{A}^{\top}+\hat{\lambda}I_{p})(\hat{\bar{K}}_{2}\otimes\hat{\bar{K}}_{1})^{\top}.$
@@
(30) $\begin{split}\bar{K}_{1}^{(t)}&:=\operatorname{argmin}_{\bar{K}_{1}\in\mathbb{P}(\mathcal{M}_{1})}\ell(\bar{K}_{1},\bar{K}_{2}^{(t-1)},\nu^{(t-1)},A^{(t-1)},\lambda^{(t-1)}),\\ \bar{K}_{2}^{(t)}&:=\operatorname{argmin}_{\bar{K}_{2}\in\mathbb{P}(\mathcal{M}_{2})}\ell(\bar{K}_{1}^{(t)},\bar{K}_{2},\nu^{(t-1)},A^{(t-1)},\lambda^{(t-1)}),\\ \nu^{(t)}&:=\operatorname{argmin}_{\nu\in\real_{+}}\ell(\bar{K}_{1}^{(t)},\bar{K}_{2}^{(t)},\nu,A^{(t-1)},\lambda^{(t-1)}),\\ A^{(t)}&:=\operatorname{argmin}_{A\in\mathcal{C}_{p_{1},p_{2},r}}\ell(\bar{K}_{1}^{(t)},\bar{K}_{2}^{(t)},\nu^{(t)},A,\lambda^{(t-1)}),\\ \lambda^{(t)}&:=\operatorname{argmin}_{\lambda\in(0,1)}\ell(\bar{K}_{1}^{(t)},\bar{K}_{2}^{(t)},\nu^{(t)},A^{(t)},\lambda).\end{split}$
@@
(31) $\operatorname{Hess}\ell(\theta^{(t-1)})[V]=-\operatorname{grad}\ell(\theta^{(t-1)}),$
@@
(34) $\nu^{(t)}=\sqrt{\frac{\textsf{tr}\left((\bar{K}^{(t)})^{-1}S(\bar{K}^{(t)})^{-\top}((1-\lambda^{(t-1)})A^{(t-1)}(A^{(t-1)})^{\top}+\lambda^{(t-1)}I_{p})^{-1}\right)}{p}},$
@@
(35) $\bar{K}_{i}^{(0)}$ $=\tilde{K}_{i}/|\tilde{K}_{i}|^{1/p_{i}},\,\nu^{(0)}=\prod_{i=1}^{2}|\tilde{K}_{i}|^{1/p_{i}},\,\lambda^{(0)}=\frac{p-\sum_{i=1}^{r}\lambda_{i}(\tilde{C}_{r})}{p-r},\,A^{(0)}=U_{r}\Lambda_{r},$
@@
Input: $\epsilon>0:$ tolerance parameter, $T\in\mathbb{N}:$ maximum number of iterations, $Y_{1},\ldots,Y_{n}\in\real^{p_{1}\times p_{2}}$: $n$ data matrices, $r\in\mathbb{N}:$ partial-isotropy rank.
Output: $(\bar{K}_{1}^{(t)},\bar{K}_{2}^{(t)},\nu^{(t)},A^{(t)},\lambda^{(t)})$.
@@
7 Numerical illustration of PICSE
We numerically illustrate the effectiveness of PICSE when the population covariance exhibits a partial isotropy core structure. [^1] Replication code is available at https://github.com/Seungbongjung/riemmCore. We randomly generate the data $Y_{1},\ldots,Y_{n}$ according to the partial isotropy core covariance model in (27), where we take $K^{1/2}=K_{2}\otimes K_{1}$ as a symmetric square root for convenience. Here we take $\lambda=0.4$, $r=4$, $(p_{1},p_{2})=(12,8),(16,10)$, and $n=p/8,p/4,p/2,p,2p$. Also, the parameter $(K_{1},K_{2},A)$ is randomly generated. We assume that $r$ is known. In practice, the value of $r$ can be determined by estimating the number of the spiked eigenvalues of the true core using the sample core in view of Kronecker-invariance [sung2025], assuming the constant non-spiked eigenvalues, e.g., [passemier2012, passemier2014].

### v3

@@ 5.2 Riemannian gradient and Hessian operator on $\mathcal{C}_{p_{1},p_{2}}^{++}$
Endow $\mathcal{C}_{p_{1},p_{2}}^{++}$ with the Euclidean metric $g^{E}$. By Proposition 3.6, the form of the tangent space $T_{C}\mathcal{C}_{p_{1},p_{2}}^{++}$ does not depend on $C\in\mathcal{C}_{p_{1},p_{2}}^{++}$. The same holds for the form of the metric, i.e., $g_{C}^{E}(U,V)=\textsf{tr}\left(U^{\top}V\right)$ for $U,V\in T_{C}\mathcal{C}_{p_{1},p_{2}}^{++}$. Thus, letting $\mathcal{W}:=T_{C}\mathcal{C}_{p_{1},p_{2}}^{++}$, which does not depend on $C$, $\mathcal{W}$ is a linear subspace of $\mathcal{S}_{p}$. Also, with any fixed basis of $\mathcal{W}$ as a coordinate on each tangent space, we have that the Christoffel symbols (see $(5.10)$ of [39]) vanish on that coordinate. Thus, $(\mathcal{C}_{p_{1},p_{2}}^{++},g^{E})$ has a zero-sectional curvature, and so $(\mathcal{C}_{p_{1},p_{2}}^{++},g^{E})$ is flat ([39], Theorem $7.10$). Now we derive the Riemannian gradient and Hessian operator on $(\mathcal{C}_{p_{1},p_{2}}^{++},g^{E})$. For a scalar-valued smooth map $f$ on $\mathcal{C}_{p_{1},p_{2}}^{++}$, denote the Euclidean derivative and Hessian operator of $f$ by $\nabla f(C)$ and $\nabla^{2}f(C)[V]$ for $C\in\mathcal{C}_{p_{1},p_{2}}^{++}$ and $V\in T_{C}\mathcal{C}_{p_{1},p_{2}}^{++}$.
@@
$\mathcal{G}(V):=V-\frac{1}{p_{2}}(I_{p_{2}}\otimes\textsf{tr}_{1}(V))-\frac{1}{p_{1}}(\textsf{tr}_{2}(V)\otimes I_{p_{1}})+\frac{\textsf{tr}(V)}{p}I_{p}$ (25)
@@
Proof.
See Appendix A.6 for the proof. ∎
Proposition 5.5.
@@
Proof.
See Appendix A.6 for the proof. ∎
@@
We derive the Riemannian gradient and Hessian operator on $(\mathcal{C}_{p_{1},p_{2},r},g^{E})$ and then deduce those on $(\mathcal{C}_{p_{1},p_{2},r}/\mathcal{O}_{r},g^{E,0})$ via quotient geometry. Here $g^{E,0}$ is the quotient metric induced by $g^{E}$. Throughout this section, we denote the Moore-Penrose pseudoinverse of a matrix $M$ by $M^{\dagger}$ (see p.50–51 of [57]). We establish the Riemannian gradient and Hessian operator on $(\mathcal{C}_{p_{1},p_{2},r},g^{E})$ as follows.
@@
Proof.
See Appendix A.7 for the proof. ∎
Proposition 5.7.
@@
Proof.
See Appendix A.7 for the proof. ∎
The Riemannian gradient and Hessian operator on $(\mathcal{C}_{p_{1},p_{2},r}/\mathcal{O}_{r},g^{E,0})$ can be derived using the results of [1, 45] (see Section $5.1$ and $5.6$ of [15] for example). For $A\in\mathcal{C}_{p_{1},p_{2},r}$, the vertical and horizontal spaces at $A$ are given by
@@
Note that $T_{A}\mathcal{C}_{p_{1},p_{2},r}=\mathcal{V}_{A}\oplus\mathcal{H}_{A}$. Then we introduce the operators on $T_{A}\mathcal{C}_{p_{1},p_{2},r}$ by [45] as
$P_{A}^{v}(W):=A\mathbf{T}_{A^{\top}A}^{-1}(2\textsf{skew}(A^{\top}W)),\quad P_{A}^{h}(W):=W-P_{A}^{v}(W).$ (26)
for $W\in T_{A}\mathcal{C}_{p_{1},p_{2},r}$. Here the operator $\mathbf{T}_{E}^{-1}(\cdot)$ is the inverse of the map $\mathbf{T}_{E}:Y\in\real^{r\times r}\rightarrow YE+EY\in\real^{r\times r}$. If $E\in\mathcal{S}_{r}^{++}$, $\mathbf{T}_{E}$ is invertible and the value of its inverse $\mathbf{T}_{E}^{-1}(V)$ for given $V$ is a unique solution to the Sylvester equation $YE+EY=V$ ([45], Lemma A.10). Also, by Section $5.3.4$ of [1], the Riemannian connection $\nabla$ on $\mathcal{C}_{p_{1},p_{2},r}/\mathcal{O}_{r}$ satisfies that
@@
for any $A\in\mathcal{C}_{p_{1},p_{2},r}$, vector fields $\eta,\xi$ on $\mathcal{C}_{p_{1},p_{2},r}/\mathcal{O}_{r}$ and the operator $P_{A}^{h}$ defined in (26). Note that $\eta^{\#}$ and $\xi^{\#}$ are horizontal lifts of $\eta$ and $\xi$, respectively. As a direct consequence of Section $3.6.2$ and $5$ of [1] and Proposition A.14 of [45], we have the following results.
Proposition 5.8.
@@ 6 Partial isotropy core shrinkage estimator
Using the geometry of $\mathcal{C}_{p_{1},p_{2},r}$, we shall propose a shrinkage estimator that shrinks the low-dimensional core toward the trivial core, $I_{p}$. Suppose $Y_{1},\ldots,Y_{n}\overset{i.i.d.}{\sim}N_{p_{1}\times p_{2}}(0,\Sigma)$ and let $K^{1/2}CK^{1/2,\top}$ be a KCD of $\Sigma$. In the estimation of $\Sigma$, note that the dimension of the space where $K=k(\Sigma)$ is living is $O(p_{1}^{2}+p_{2}^{2})=o(p^{2})$, whereas that of the space where $C=c(\Sigma)$ is living is $O(p^{2})$ by Theorem 3.5. Thus, the difficulty of the estimation arises mainly from estimating $C$, particularly in a high-dimensional regime where $p>n$.
As a remedy, we consider a partial-isotropy structure on $C$ to introduce a low-dimensional structure to $C$ as discussed in Section 1 and 2.1. By Proposition 2.2, $C=(1-\lambda)AA^{\top}+\lambda I_{p}$ for some $\lambda\in(0,1)$ and $A\in\mathcal{C}_{p_{1},p_{2},r}$, with $r>p_{1}/p_{2}+p_{2}/p_{1}$. Thus, $\Sigma=K^{1/2}((1-\lambda)AA^{\top}+\lambda I_{p})K^{1/2,\top}$, leading to the partial-isotropy core covariance model in (2). Namely,
$Y_{1},\ldots,Y_{n}\overset{i.i.d.}{\sim}N_{p_{1}\times p_{2}}(0,K^{1/2}((1-\lambda)AA^{\top}+\lambda I_{p})K^{1/2,\top}).$ (27)
@@
Note that $\lambda$ denotes the shrinkage amount of the low-dimensional covariance $K^{1/2}AA^{\top}K^{1/2,\top}$ toward the separable part $K:=K^{1/2}K^{1/2,\top}$. Hence, $\lambda$ quantifies an effective departure from the separability assumption on $\Sigma$, i.e., $\Sigma=K$, where the correlation structure of $\Sigma$ may be too simplified as $p$ grows.
For the identifiability of the covariance model above, recall that there is an ambiguity in identifying the factors of $K^{1/2}:=K_{2}\otimes K_{1}$ due to a constant factor. To avoid this ambiguity, we shall reparameterize $K^{1/2}$ by $K^{1/2}:=\nu(\bar{K}_{2}\otimes\bar{K}_{1})$, where $\bar{K}_{i}\in\mathbb{P}(\mathcal{S}_{p_{i}}^{++})$ or $\mathbb{P}(\mathcal{L}_{p_{i}}^{++})$, and $\nu>0$. Under this parameterization, the parameters constituting the partial-isotropy core covariance are identifiable with $A$ up to right-rotation.
Proposition 6.1.
Given $p_{1}/p_{2}+p_{2}/p_{1}<r<p$, define the parameter space
$\Theta:=\mathbb{P}(\mathcal{M}_{1})\times\mathbb{P}(\mathcal{M}_{2})\times\real_{+}\times\mathcal{C}_{p_{1},p_{2},r}\times(0,1),$
where $(\mathcal{M}_{1},\mathcal{M}_{2})$ is either $(\mathcal{L}_{p_{1}}^{++},\mathcal{L}_{p_{2}}^{++})$ or $(\mathcal{S}_{p_{1}}^{++},\mathcal{S}_{p_{2}}^{++})$. Suppose a smooth map $\Omega:\Theta\rightarrow\mathcal{S}_{p}^{++}$ is defined by
$\Omega(\tau)=K^{1/2}((1-\lambda)AA^{\top}+\lambda I_{p})K^{1/2,\top},$
where $K^{1/2}=\nu(\bar{K}_{2}\otimes\bar{K}_{1})$ and $\tau=(\bar{K}_{1},\bar{K}_{2},\nu,A,\lambda)$. For $\tau_{i}=(\bar{K}_{1}^{i},\bar{K}_{2}^{i},\nu^{i},A^{i},\lambda^{i})\in\Theta$ with $i=1,2$, $\Omega(\tau_{1})=\Omega(\tau_{2})$ if and only if for some $O\in\mathcal{O}_{r}$,
$\bar{K}_{1}^{1}=\bar{K}_{1}^{2},\quad\bar{K}_{2}^{1}=\bar{K}_{2}^{2},\quad\nu^{1}=\nu^{2},\quad A^{1}=A^{2}O,\quad\lambda^{1}=\lambda^{2}.$
Proof.
See Appendix A.7 for the proof. ∎
@@
Now we propose a partial-isotropy core shrinkage estimator (PICSE) by $\hat{K}^{1/2}((1-\hat{\lambda})\hat{A}\hat{A}^{\top}+\hat{\lambda})\hat{K}^{1/2,\top}$, where $\hat{\theta}$ is a MLE of the parameter $\theta$. We shall consider both square roots $K^{1/2}\in\mathcal{S}_{p_{1},p_{2}}^{++}$ or $\mathcal{L}_{p_{1},p_{2}}^{++}$ in the estimation. Given the data $Y_{1},\ldots,Y_{n}$ according to the model in (27), the negative log-likelihood is given by
$\begin{split}\ell(\bar{K}_{1},\bar{K}_{2},\nu,A,\lambda)&:=\textsf{tr}\left(\bar{K}^{-1/2}S\bar{K}^{-1/2,\top}((1-\lambda)AA^{\top}+\lambda I_{p})^{-1}\right)/\nu^{2}\\ &+\log|(1-\lambda)AA^{\top}+\lambda I_{p}|+2p\log\nu.\end{split}$ (28)
@@
$\hat{\Sigma}_{\texttt{PICSE}}:=\hat{\nu}^{2}(\hat{\bar{K}}_{2}\otimes\hat{\bar{K}}_{1})((1-\hat{\lambda})\hat{A}\hat{A}^{\top}+\hat{\lambda}I_{p})(\hat{\bar{K}}_{2}\otimes\hat{\bar{K}}_{1})^{\top}.$ (29)
@@
$\begin{split}\bar{K}_{1}^{(t)}&:=\operatorname{argmin}_{\bar{K}_{1}\in\mathbb{P}(\mathcal{M}_{1})}\ell(\bar{K}_{1},\bar{K}_{2}^{(t-1)},\nu^{(t-1)},A^{(t-1)},\lambda^{(t-1)}),\\ \bar{K}_{2}^{(t)}&:=\operatorname{argmin}_{\bar{K}_{2}\in\mathbb{P}(\mathcal{M}_{2})}\ell(\bar{K}_{1}^{(t)},\bar{K}_{2},\nu^{(t-1)},A^{(t-1)},\lambda^{(t-1)}),\\ \nu^{(t)}&:=\operatorname{argmin}_{\nu\in\real_{+}}\ell(\bar{K}_{1}^{(t)},\bar{K}_{2}^{(t)},\nu,A^{(t-1)},\lambda^{(t-1)}),\\ A^{(t)}&:=\operatorname{argmin}_{A\in\mathcal{C}_{p_{1},p_{2},r}}\ell(\bar{K}_{1}^{(t)},\bar{K}_{2}^{(t)},\nu^{(t)},A,\lambda^{(t-1)}),\\ \lambda^{(t)}&:=\operatorname{argmin}_{\lambda\in(0,1)}\ell(\bar{K}_{1}^{(t)},\bar{K}_{2}^{(t)},\nu^{(t)},A^{(t)},\lambda).\end{split}$ (30)
@@
$\operatorname{Hess}\ell(\theta^{(t-1)})[V]=-\operatorname{grad}\ell(\theta^{(t-1)}),$ (31)
@@
$\nu^{(t)}=\sqrt{\frac{\textsf{tr}\left((\bar{K}^{(t)})^{-1}S(\bar{K}^{(t)})^{-\top}((1-\lambda^{(t-1)})A^{(t-1)}(A^{(t-1)})^{\top}+\lambda^{(t-1)}I_{p})^{-1}\right)}{p}},$ (34)
@@
$\bar{K}_{i}^{(0)}$ $=\tilde{K}_{i}/|\tilde{K}_{i}|^{1/p_{i}},\,\nu^{(0)}=\prod_{i=1}^{2}|\tilde{K}_{i}|^{1/p_{i}},\,\lambda^{(0)}=\frac{p-\sum_{i=1}^{r}\lambda_{i}(\tilde{C}_{r})}{p-r},\,A^{(0)}=U_{r}\Lambda_{r},$ (35)
@@
$\epsilon>0:$ tolerance parameter, $T\in\mathbb{N}:$ maximum number of iterations, $Y_{1},\ldots,Y_{n}\in\real^{p_{1}\times p_{2}}$: $n$ data matrices, $r\in\mathbb{N}:$ partial-isotropy rank.
@@
7 Illustration of PICSE
We illustrate the effectiveness of PICSE based on synthetic data. [^1] Replication code is available at https://github.com/Seungbongjung/riemmCore. We randomly generate the random matrices $Y_{1},\ldots,Y_{n}$ according to $N_{p_{1}\times p_{2}}(0,\Sigma)$, where $\Sigma$ is given as follows; for $K^{1/2}=K_{2}\otimes K_{1}\in\mathcal{S}_{p_{1},p_{2}}^{++}$, $A\in\mathcal{C}_{p_{1},p_{2},r}$, $\lambda\in(0,1)$, and a diagonal $D\in\mathcal{C}_{p_{1},p_{2}}^{++}$,
  • (M1) $\Sigma=K^{1/2}((1-\lambda)AA^{\top}+\lambda I_{p})K^{1/2,\top}$,
  • (M2) $\Sigma=K^{1/2}((1-\lambda)AA^{\top}+\lambda D)K^{1/2,\top}$.
The model (M1) is the partial isotropy core covariance model in (27), and thus $c(\Sigma)=(1-\lambda)AA^{\top}+\lambda I_{p}$. On the other hand, the model (M2) is a variant of (M1), motivated by a general factor covariance model [42, 10]. We also consider this model to examine the robustness of PICSE under a broader class of covariance models for matrix-variate data, containing the partial isotropy core covariance model. Under this model, $c(\Sigma)=(1-\lambda)AA^{\top}+\lambda D$. By the linear system in (4) that defines the core, if a diagonal $D$ is a core, $(1-\lambda)AA^{\top}+\lambda D$ is again a core. One can easily generate such a $D$ by randomly generating a positive definite diagonal $\tilde{D}$ and then taking its core D=c(D~)D=c($\tilde${D}). This is because the Kronecker MLE of any positive definite diagonal matrix $\tilde{D}$ is again diagonal, and so is its core ([33], Corollary $1$).
To investigate how PICSE behaves by varying degrees of how $\Sigma$ is separable, we take $\lambda=0.2,0.4,0.6,0.8$. Note that the smaller $\lambda$ is, the less separable $\Sigma$ is. Also, we consider $r=3,5$, $(p_{1},p_{2})=(16,12),(18,8)$, and $n=p/8,p/4,p/2,p,2p$. The other parameters $(K_{1},K_{2},A,D)$ are randomly generated. We assume the known $r$. In practice, the value of $r$ can be determined by estimating the number of the spiked eigenvalues of the true core using the sample core in view of Kronecker-invariance [63], assuming the constant non-spiked eigenvalues, e.g., [52, 53].